### v2

# Continuous-time reinforcement learning for optimal switching over multiple regimes

Abstract
This paper studies the continuous-time reinforcement learning (RL) for optimal switching problems across multiple regimes. We consider a type of exploratory formulation under entropy regularization where the agent randomizes both the timing of switches and the selection of regimes through the generator matrix of an associated continuous-time finite-state Markov chain. We establish the well-posedness of the associated system of Hamilton-Jacobi-Bellman (HJB) equations and provide a characterization of the optimal policy. The policy improvement and the convergence of the policy iterations are rigorously established by analyzing the system of equations. We also show the convergence of the value function in the exploratory formulation towards the value function in the classical formulation as the temperature parameter vanishes. Finally, a reinforcement learning algorithm is devised and implemented by invoking the policy evaluation based on the martingale characterization. Our numerical examples with the aid of neural networks illustrate the effectiveness of the proposed RL algorithm.
Keywords: Optimal regime switching, multiple regimes, continuous-time reinforcement learning, system of HJB equations, policy improvement, policy iteration convergence
1 Introduction
The optimal switching problem across multiple regimes entails solving a stochastic optimization problem in which the admissible strategies are formalized by sequences of discrete interventions. A decision-maker in this context faces two basic questions: (i) when to switch from the current regime to another, and (ii) which regime to select when the decision of switching is made. These problems are characterized by their hybrid nature, combining continuous state dynamics with discrete control actions, where each switch between regimes typically incurs a cost while different regimes yield different reward outcomes. Over recent decades, the optimal switching problem has found extensive applications across different fields. Seminal work includes carmona2008pricing on pricing asset scheduling, carmona2010valuation on energy storage valuation, porchet2009valuation on power plant valuation, and olofsson2022management on hydropower production planning, among others.
The classical stochastic control problem typically assumes a fully known and accurate underlying model. However, this assumption of complete model knowledge often turns out to be unrealistic in practical applications. RL offers a powerful framework for learning optimal strategies in the unknown environment through trial-and-error interactions. While most conventional RL algorithms are designed in discrete-time settings, many real-world applications evolve continuously in time, motivating a systemic study in theories and algorithms for the continuous-time RL approach. Within the continuous-time framework, decision-makers face the fundamental exploration-exploitation trade-off in a continuous-time manner: whether to exploit current knowledge by executing the best-known policy or to explore alternative actions to gather information for potential long-term improvement. wang2020reinforcement addressed this problem by introducing an entropy-regularization on the randomized policy to encourage the exploration. This fundamental study spurred further pioneer studies of theories and algorithms in the continuous-time exploratory framework including jia2022policy; jia2022gradient; jia2023q, laying the foundations for the policy evaluation, the policy gradient, and the continuous-time q-learning, respectively. Later, the well-posedness of the exploratory HJB equation, the convergence of policy iterations and the regret analysis have also been examined in tang2022exploratory; huang2025convergence; tran2025policy; t-z-regret.
In addition, vast extensions and applications of continuous-time RL algorithms in various context have been considered in the recent literature. To name a few, wu2024reinforcement addressed the continuous-time mean-variance portfolio selection problem in regime-switching markets with unobservable states using reinforcement learning approach; bo2025optimal extended the q-learning theory in the model of reflected diffusion processes and applied it to learn the optimal tracking portfolio in incomplete markets; wei2025continuous generalized the continuous-time q-learning to mean-field control problems within McKean-Vlasov diffusion models; wyy2024 further developed the continuous-time q-learning for both mean-field control and mean-field game problems from the perspective of the representative agent; gao2024reinforcement studied the extension of q-learning in jump-diffusion models; bo2024continuous examined the same jump-diffusion model by invoking the Tsallis entropy; dong2024randomized investigated the optimal stopping in an exploratory framework by considering the randomization of stopping time via the intensity control; dianetti2024exploratory utilized the randomization of stopping times as singular control and studied its exploratory formulation under residual entropy regularization; dai2024learning exploited the penalization method to transform the optimal stopping problem to an optimal control problem for which the entropy regularization is formalized; liang2025reinforcement proposed a continuous-time RL framework for singular stochastic control problems without entropy regularization, characterizing the optimal control through singular control laws; liang2025reinforcement-2 further proposed a type of randomization of the singular control laws in liang2025reinforcement by considering an auxiliary singular control and entropy regularization, which lead to a time-inconsistent two-stage optimal control problem such that the task is to learn the time-consistent equilibrium.
Despite these advancements of continuous-time RL in different model setups, its application to optimal regime switching problems remains relatively underexplored. This paper studies the exploratory formulation of the optimal regime switching with multiple regimes and bridges its connection to the classical optimal switching problem as the entropy regularization vanishes. To this end, we propose a type of exploratory formulation where the decision-maker randomizes both switching time and the selection of the targeted regime state by invoking a generator matrix of an associated continuous-time Markov chain (CTMC) defined on finite state space. The entropy regularization on the generator is imposed to encourage the exploration. Specifically, we utilize the inherent property of the CTMC—particularly its jump times and state transitions—to determine the switching decision. This formulation, governed by the control of the CTMC’s generator matrix, transformed the randomized switching problem into an optimal control problem.
We summarize the main contributions of the present paper as follows:
  • (i)
    We derive the system of exploratory HJB equations and establish the existence of a bounded classical solution to this system (see Lemma 3.2) by resorting to some established partial differential equation (PDE) theories together with a tailor-made truncation argument. Furthermore, we prove its uniqueness and demonstrate through a verification theorem (see Proposition 3.3) that this solution coincides with the value function.
  • (ii)
    We employ the policy iteration (PI) method to learn the optimal strategy through iterative updates and prove the policy improvement result in Proposition 4.1. As the main result of this paper, in the context of PDE system, we establish the convergence result of the policy iteration in Theorem 4.2 with an explicit convergence rate, which is new to the literature.
  • (iii)
    We also draw the connection to the classical optimal switching problem by establishing the convergence of the value function in the exploratory formulation towards the value function of the classical optimal switching problem as the temperature parameter approaches zero. To this end, we resort to some delicate stability analysis of viscosity solutions of the PDE system, see Lemma 4.3 and Theorem 4.4. In particular, it is shown that the solution of the system of PDEs will converge to the solution of the system of variational inequalities as the temperature parameter tends to zero.
  • (iv)
    We develop a reinforcement learning algorithm by implementing a policy evaluation method based on martingale characterization, which calls for the stochastic approximation when using the martingale orthogonality condition. We obtain an explicit error analysis for the convergence of this stochastic approximation method in Theorem 5.4. To illustrate the effectiveness of our proposed RL algorithm, we conducted numerical experiments in two examples with satisfactory iteration convergence, both necessitate the application of neural networks to parameterize the targeted functions.
Let us also briefly compare the present work with three recent related studies. denkert2025control introduced a control randomization method without entropy regularization in continuous-time RL with the application to optimal switching problems. They developed an Actor-Critic policy gradient algorithm that alternately learns the value function and the optimal intensity policy. In contrast, our paper propose a different randomization approach for the optimal switching problem, utilizing the generator matrix of a CTMC and incorporating entropy regularization to encourage the exploration. A key advantage of our formulation is that the optimal policy depends explicitly on the value function itself, without requiring any of its derivatives. This allows us to parameterize both the policy and the value function using the same set of parameters. More recently, dai2025reinforcement developed a RL approach to identify arbitrage strategies in stock index futures. Following the randomization method in dong2024randomized, they randomized the switching times in dai2025reinforcement using the Cox processes and formulated the problem as an optimal switching problem with three regimes where the state process is independent of the regimes. In comparison, we consider an exploratory framework for a more general multi-regime optimal switching problem, where the state process dynamics can also depend on the regime states. Furthermore, we rigorously establish the convergence of the policy iterations with an explicit convergence rate and also show the convergence as the entropy regularization vanishes. Finally, our work differs from cao2025two, which studied a randomization scheme for impulse control problems characterized by fixed points of compound operators combining regularized nonlocal and stopping operators. In contrast, our distinct exploratory formulation leads to the study of PDE system instead of a single PDE problem, for which we need to develop some delicate analysis for the system of equations to deduce some desired convergence results.
The remainder of this paper is organized as follows. Section 2 reviews the classical optimal switching problem and presents preliminary results on viscosity solutions to the associated system of HJB variational inequalities. Section 3 introduces the exploratory formulation of the optimal switching problem, providing a regularity analysis of the value function and the characterization of the optimal policy. Section 4 establishes both policy improvement and the convergence result of the policy iteration. Moreover, the convergence behavior of the exploratory solution as the temperature parameter vanishes is also discusses therein. Section 5 develops a reinforcement learning algorithm that implements the martingale-based policy evaluation and the previous policy iteration, accompanied by an error analysis for the proposed algorithm. Finally, Section 6 presents some numerical examples demonstrating the satisfactory performance of our proposed RL algorithm.
Notations. We specify the following list of notations for the rest of this paper.
  • $\mathds{R}^{n}$ denotes the $n$-dimensional Euclidean space. For all $x=(x_{1},\cdots,x_{n}),y=(y_{1},\cdots,y_{n})\in\mathds{R}^{n}$, we denote by $\cdot$ the scalar product and by $|\cdot|$ the Euclidean norm:
    $x\cdot y=\sum_{i=1}^{n}x_{i}y_{i},\quad|x|=\sqrt{x\cdot x}=\sqrt{\sum_{i=1}^{n}x_{i}^{2}}.$
  • $\mathds{R}^{n\times d}$ is the set of real-valued $n\times d$ matrices. For $\sigma\in\mathds{R}^{n\times d}$, we denote by $\sigma^{\top}$ the transpose matrix of $\sigma$. For $A=(a_{ij})_{1\leq i,j\leq n}\in\mathds{R}^{n\times n}$, $\text{tr}(A)=\sum_{i=1}^{n}a_{ii}$ is the trace of $A$. We define the matrix norm on $\mathds{R}^{n\times d}$ as $|\sigma|=(\text{tr}(\sigma\sigma^{\top}))^{\frac{1}{2}}$.
  • For ${\cal O}\subset\mathds{R}^{n}$, $C^{k}({\cal O})$ is the space of all real-valued continuous functions on $\cal O$} with continuous derivatives up to order $k$. For $T\geq 0$, $C^{1,2}([0,T]\times{\cal O})$ is the space of real-valued functions $u$ on $0,T]\times{\cal O}$ whose partial derivatives $\frac{\partial u}{\partial t},\frac{\partial u}{\partial x_{i}},\frac{\partial^{2}u}{\partial x_{i}x_{j}}$, $1\leq i,j\leq n$, exist and are continuous on $0,T]\times{\cal O}$. For $u\in C^{2}({\cal O})$, we denote by $D_{x}u$ the gradient vector of $u$ and $D_{x}^{2}u$ the Hessian matrix of $u$.
  • For points $P=(t,x),P^{\prime}=(t,x)\in[0,T]\times\mathds{R}^{n}$, we define the parabolic distance between $P$ and $P^{\prime}$ by
    $d(P,P^{\prime})=(|t-t^{\prime}|+|x-x^{\prime}|^{2})^{\frac{1}{2}}.$
  • For ${\cal D}\subset[0,T]\times\mathds{R}^{n}$ and $\alpha\in(0,1)$ we introduce the following norms for functions defined on $\cal D$}:
[…]
    We shall say that function $u(t,x)$ is in $C^{q}({\cal D})$ if $u||_{C^{q}({\cal D})}$ is finite ($q=0,\alpha,1+\alpha,2+\alpha$).
2 Classical Optimal Switching Problem
This section first reviews the classical optimal switching problem and introduce some preliminary results on viscosity solutions to the associated system of HJB variational inequalities.
We fix a complete probability space $(\Omega,\mathcal{F},\mathbb{P})$, supporting a $d$-dimensional standard Brownian motion $W=(W_{t})_{t\geq 0}$. We denote by $\mathbb{F}$ the complete and right continuous filtration generated by $W$. The terminal time is denoted by $T>0$. Let us introduce the domain ${\cal D}:=[0,T)\times\mathds{R}^{n}$, then the closure of $\cal D$} is given by $\overline{{\cal D}}=[0,T]\times\mathds{R}^{n}$.
We then define the set $\mathcal{A}_{t}$ of admissible switching controls at time $t\in[0,T]$ as the set of double sequences $\alpha=(\tau_{k},\kappa_{k})_{k\geq 0}$, where $\left(\tau_{k}\right)_{k\geq 0}$ is a non-decreasing sequence of $\mathbb{F}$-stopping times with $\tau_{0}=t$ and $\lim_{k\rightarrow\infty}\tau_{k}>T$; $\kappa_{k}$ is an $\mathcal{F}_{\tau_{k}}$-measurable random variable valued in the set $\mathbb{I}_{m}=\{1,2,\cdots,m\}$. With a strategy $\alpha=\left(\tau_{k},\kappa_{k}\right)_{k\geq 0}\in\mathcal{A}_{t}$ and an initial regime value $i\in\mathbb{I}_{m}$, we associate the process $(I_{s}^{t,i})_{s\geq t}$ defined by
$I_{s}^{t,i}=\sum_{k\geq 0}\kappa_{k}{\bf 1}_{s\in[\tau_{k},\tau_{k+1})},\penalty 10000\ s\geq t,\quad I^{t,i}_{t-}=\kappa_{0}=i.$ (2.1)
Given $(t,x,i)\in[0,T]\times\mathds{R}^{n}\times\mathbb{I}_{m}$, and a switching control $\alpha\in\mathcal{A}_{t}$, we consider the controlled diffusion $X^{t,x,i,\alpha}=(X_{s}^{t,x,i,\alpha})_{s\in[t,T]}$ governed by the SDE:
$dX^{t,x,i,\alpha}_{s}=\mu(s,X^{t,x,i,\alpha}_{s},I_{s}^{t,i})ds+\sigma(s,X^{t,x,i,\alpha}_{s},I_{s}^{t,i})dW_{s},\quad s\in(t,T],$ (2.2)
with $X_{t}^{t,x,i,\alpha}=x$. We have the following assumptions for the model coefficients.
Assumption 2.1.
  • (i)
    The drift $\mu(\cdot,\cdot,\cdot):[0,T]\times\mathds{R}^{n}\times\mathbb{I}_{m}\to\mathds{R}^{n}$ and volatility $\sigma(\cdot,\cdot,\cdot):[0,T]\times\mathds{R}^{n}\times\mathbb{I}_{m}\to\mathds{R}^{n\times d}$ are uniformly Lipschitz continuous with respect to $x$, that is, there exists a constant $L>0$ such that
    $|\mu(s,x_{1},i)-\mu(s,x_{2},i)|+|\sigma(s,x_{1},i)-\sigma(s,x_{2},i)|\leq L|x_{1}-x_{2}|$ (2.3)
    for all $(s,x_{1},x_{2},i)\in[0,T]\times\mathds{R}^{2n}\times\mathbb{I}_{m}$.
  • (ii)
    The drift $\mu$ and volatility $\sigma$ have linear growth in $x$, that is, there exists a constant $C>0$ such that
    $|\mu(s,x,i)|+|\sigma(s,x,i)|\leq C(1+|x|)$
    for all $(s,x,i)\in[0,T]\times\mathds{R}^{n}\times\mathbb{I}_{m}$.
  • (iii)
    There exist some constant $\sigma_{0}>0$ such that, for all $(t,x,i)\in\overline{{\cal D}}\times\mathbb{I}_{m}$ and $\xi\in\mathds{R}^{n}$,
    $\xi\sigma(t,x,i)\sigma^{\top}(t,x,i)\xi^{\top}\geq\sigma_{0}\xi\xi^{\top}.$
The expected total profit with the initial state $(t,x,i)$ and the impulse control $\alpha=\left(\tau_{k},\kappa_{k}\right)_{k\geq 0}\in\mathcal{A}_{t}$ is given by
$J_{i}(t,x;\alpha)=\mathbb{E}\bigg[\int_{t}^{T}f(s,X_{s}^{t,x,i,\alpha},I_{s}^{t,i})ds-\sum\limits_{k=1}^{\infty}g_{\kappa_{k-1}\kappa_{k}}{\bf 1}_{\{\tau_{k}\leq T\}}+h(X_{T}^{t,x,i,\alpha})\bigg],$ (2.4)
where $f(\cdot,\cdot,\cdot):[0,T]\times\mathds{R}^{n}\times\mathbb{I}_{m}\to\mathds{R}$ is the running profit function, $h(\cdot):\mathds{R}^{n}\to\mathds{R}$ is the terminal reward function, and the constant $g_{ij}$ denotes the cost for switching from regime $i$ to $j$ for all $i\neq j$. We also impose the following assumptions.
Assumption 2.2.
  • (i)
    For $i\in\mathbb{I}_{m}$, the running profit $f(\cdot,\cdot,i)$ and terminal reward $h(\cdot)$ are assumed to be continuous. Furthermore, there exists a constant $K_{f,h}>0$ such that
    $|f(t,x,i)|+|h(x)|\leq K_{f,h},\quad\forall(t,x,i)\in[0,T]\times\mathds{R}^{n}\times\mathbb{I}_{m}.$ (2.5)
  • (ii)
    For $i,j\in\mathbb{I}_{m}$ with $j\neq i$, the cost for switching from regime $i$ to $j$ is positive, that is, $g_{ij}>0$, with the convention $g_{ii}=0$. For $i,j,k\in\mathbb{I}_{m}$ with $j\neq i,k$, it is less expensive to switch directly in one step from regime $i$ to $k$ than in two steps via an intermediate regime $j$, that is, $g_{ik}<g_{ij}+g_{jk}$.
The objective is to maximize the expected total profit over all strategies $\alpha$. Accordingly, the classical value functions is defined by
$V_{i}(t,x)=\sup\limits_{\alpha\in\mathcal{A}_{t}}J_{i}(t,x;\alpha),\quad(t,x,i)\in[0,T]\times\mathds{R}^{n}\times\mathbb{I}_{m}.$ (2.6)
We now consider the following system of HJB variational inequalities, for $i\in\mathbb{I}_{m}$,
$\begin{cases}\displaystyle\min\left\{-\frac{\partial V_{i}(t,x)}{\partial t}-\mathcal{L}^{i}_{x}V_{i}(t,x)-f(t,x,i),V_{i}(t,x)-\max_{j\neq i}(V_{j}(t,x)-g_{ij})\right\}=0,\quad(t,x)\in{\cal D},\\ \displaystyle V_{i}(T,x)=h(x),\quad x\in\mathds{R}^{n},\end{cases}$ (2.7)
where the operator $\mathcal{L}_{x}^{i}$ with $i\in\mathbb{I}_{m}$ is defined by
$\mathcal{L}_{x}^{i}l(t,x):=\mu(t,x,i)D_{x}l(t,x)+\frac{1}{2}\text{tr}(\sigma\sigma^{\top}(t,x,i)D_{x}^{2}l(t,x)),\quad\text{for}\penalty 10000\ l(t,\cdot)\in C^{2}(\mathds{R}^{n}).$
The value function $(V_{1},\cdots,V_{m})$ can be characterized as the viscosity solution of system (2.7), which is defined as below.
Definition 2.1.
Let $(u_{1},\cdots,u_{m})$ be a $m$-uplet of functions defined on $\overline{{\cal D}}$, $\mathds{R}$-valued and such that $u_{i}(T,x)=h(x)$ for any $(i,x)\in\mathbb{I}_{m}\times\mathds{R}^{n}$. The $m$-uplet $(u_{1},\cdots,u_{m})$ is called:
  • (i)
    a viscosity supersolution (respectively, subsolution) of system (2.7) if, for each $i\in\mathbb{I}_{m}$, $u_{i}$ is lower-semicontinuous (respectively, upper-semicontinuous) on $\cal D$} and for any $(t_{0},x_{0})\in{\cal D}$ and any test function $\varphi_{i}\in C^{1,2}({\cal D})$ such that $(t_{0},x_{0})$ is a local minimum point of $u_{i}-\varphi_{i}$ (respectively, maximum), we have
    $\min\Bigg\{$ $-\frac{\partial\varphi_{i}(t_{0},x_{0})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{0},x_{0})-f(t_{0},x_{0},i),$
    $\qquad\qquad\quad u_{i}(t_{0},x_{0})-\max_{j\neq i}(u_{j}(t_{0},x_{0})-g_{ij})\Bigg\}\geq 0\penalty 10000\ \text{(respectively, $\leq 0$)};$
  • (ii)
    a viscosity solution of system (2.7) if it both a viscosity supersolution and subsolution.
By using a similar proof of Theorem 5.1 in el2013stochastic, we have the comparison principle for the system (2.7).
Lemma 2.3 (Comparison Principle).
Suppose Assumptions 2.1 and 2.2 hold. Let $(u_{1},\cdots,u_{m})$ be a bounded viscosity supersolution of system (2.7) and $(v_{1},\cdots,v_{m})$ be a bounded viscosity subsolution of system (2.7). Then $v_{i}(t,x)\leq u_{i}(t,x)$ for all $(t,x,i)\in\overline{{\cal D}}\times\mathbb{I}_{m}$.
Lemma 2.3 will help the proof of uniqueness of viscosity solution. The next result relates the value function $(V_{1},\cdots,V_{m})$ to the system of variational inequalities.
Theorem 2.4.
Under Assumptions 2.1 and 2.2, the value function $(V_{1},\cdots,V_{m})$ given by (2.6) is the unique bounded viscosity solution of system (2.7).
Proof.
We begin by proving that the value function $(V_{1},\cdots,V_{m})$ defined by (2.6) is bounded. By Assumption 2.2, for any $(i,t,x)\in\mathbb{I}_{m}\times\overline{{\cal D}}$ and $\alpha\in{\cal A}_{t}$,
$J_{i}(t,x,\alpha)$ $\leq\mathbb{E}\bigg[\int_{t}^{T}f(s,X_{s}^{t,x,i,\alpha},I_{s}^{t,i})ds+h(X_{T}^{t,x,i,\alpha})\bigg]$
$\leq(T-t)K_{f,h}+K_{f,h},$
which implies $V_{i}(t,x)\leq(T-t)K_{f,h}+K_{f,h}$. For the lower bound, consider the no-switching control $\tau_{n}=\infty$, $n\geq 1$, i.e., $I_{s}^{t,i}=i$, $s\geq t$. Applying Assumption 2.2 again yields
$V_{i}(t,x)$ $\geq\mathbb{E}\bigg[\int_{t}^{T}f(s,X_{s}^{t,x,i,\alpha},I_{s}^{t,i})ds+h(X_{T}^{t,x,i,\alpha})\bigg]$
$\geq-(T-t)K_{f,h}-K_{f,h}.$
Therefore, the value function is bounded. As it is bounded, it follows from Proposition 4.2 in bouchard2009stochastic and Lemma 2.3 that the value function $(V_{1},\cdots,V_{m})$ is the unique bounded viscosity solution of system (2.7). ∎
3 Exploratory Formulation under Entropy Regularization
In this section, we introduce our exploratory formulation of the optimal switching problem, and study the well-posedness of the associated exploratory HJB system as well as the verification theorem.
To explore the system and reward, we let the agent randomize the choice of the stopping times and the regimes that she would like to switch to. Let $I:=(I_{t})_{t\geq 0}$ denote a continuous-time finite-state Markov chain with state space $\mathbb{I}_{m}$, which is independent of the Brownian motion $W$. The randomization is achieved by considering the choice of the generator, ${\bm{\pi}}=(\pi^{ij}_{t})_{i,j\in\mathbb{I}_{m},t\in[0,T]}$, of the Markov chain $I$. For $i\neq j$, $\pi_{t}^{ij}$ is the instantaneous intensity of the transition of $I$ from state $i$ to state $j$ at time $t$. Here, for each $t\in[0,T]$, $\pi^{ij}_{t}\geq 0$, for $i\neq j$ and $\sum_{j=1}^{m}\pi_{t}^{ij}=0$.
Given $(t,x,i)\in[0,T]\times\mathds{R}^{n}\times\mathbb{I}_{m}$, we consider the controlled diffusion $X=(X_{s})_{s\in[t,T]}$ defined by the following SDE:
$dX_{s}=\mu(s,X_{s},I_{s})ds+\sigma(s,X_{s},I_{s})dW_{s},\quad s\in(t,T].$ (3.1)
with $X_{t}=x$ and $I_{t}=i$. For $k\geq 1$, denote by $\tau_{k}$ the $k$-th jump time of process $I$ with $\tau_{0}=0$ and $\kappa_{k}:=I_{\tau_{k}}$. For $t\geq 0$, let $\mathbb{U}_{t}$ be the set of all admissable policies $(\pi^{ij})_{i,j\in\mathbb{I}_{m}}$ such that for every $i,j\in\mathbb{I}_{m}$, the process $\pi^{ij}=(\pi^{ij}_{s})_{s\in[t,T]}$ is $\mathbb{F}$-adapted and satisfies (i) for $i\neq j$, $\pi^{ij}_{s}\geq 0$ for all $s\in[t,T]$; (ii) for every $i\in\mathbb{I}_{m}$, $\sum_{j=1}^{m}\pi_{s}^{ij}=0$, for all $s\in[t,T]$.
For ${\bm{\pi}}\in\mathbb{U}_{t}$, denote by ${\bm{\pi}}^{i}=(\pi^{ij})_{j\in\mathbb{I}_{m}}$ for $i\in\mathbb{I}_{m}$. To encourage the exploration, we adopt the normalized entropy similar to dong2024randomized that $R({\bm{\pi}},i):=\sum_{j\neq i}\pi^{ij}-\pi^{ij}\log\pi^{ij}$ for $i\in\mathbb{I}_{m}$. The exploratory formulation of objective functional under entropy regularizer is given by, for $(t,x,i)\in[0,T]\times\mathds{R}^{n}\times\mathbb{I}_{m}$ and ${\bm{\pi}}=(\pi^{ij}_{s})_{i,j\in\mathbb{I}_{m},s\in[t,T]}\in\mathbb{U}_{t}$,
$J_{i}^{\lambda}(t,x;{\bm{\pi}}):=\mathbb{E}_{t,x,i}\bigg[\int^{T}_{t}f(s,X_{s},I_{s})ds-\sum_{k=1}^{\infty}g_{\kappa_{k-1}\kappa_{k}}{\bf 1}_{\{\tau_{k}\leq T\}}+\lambda\int^{T}_{t}R({\bm{\pi}}_{s},I_{s})ds+h(X_{T})\bigg],$ (3.2)
where $\mathbb{E}_{t,x,i}[\cdot]:=\mathbb{E}[\cdot|X_{t}=x,I_{t}=i]$, and $\lambda>0$ is the temperature parameter. Furthermore, the optimal value function is denoted by
$V_{i}^{\lambda}(t,x)=\sup_{{\bm{\pi}}\in\mathbb{U}_{t}}J_{i}^{\lambda}(t,x;{\bm{\pi}}).$ (3.3)
Applying the dynamic programming arguments (c.f. Section 5.3.2 in pham2009continuous), we derive the system of coupled HJB equations as follows: for $i\in\mathbb{I}_{m}$,
$\begin{cases}\displaystyle\frac{\partial V^{\lambda}_{i}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}V^{\lambda}_{i}(t,x)+f(t,x,i)\\ \displaystyle\qquad\quad+\sup_{{\bm{\pi}}_{i}}\left\{\sum_{j\neq i}\pi_{ij}(V^{\lambda}_{j}(t,x)-g_{ij}-V^{\lambda}_{i}(t,x))+\lambda\sum_{j\neq i}(\pi_{ij}-\pi_{ij}\log\pi_{ij})\right\}=0,\penalty 10000\ (t,x)\in{\cal D},\\ \displaystyle V^{\lambda}_{i}(T,x)=h(x),\penalty 10000\ x\in\mathds{R}^{n}.\end{cases}$ (3.4)
Using the first-order condition, we arrive at the characterization of the optimal feedback policy by
$\pi_{ij}^{*}(t,x)=\exp\left(\frac{V^{\lambda}_{j}(t,x)-g_{ij}-V^{\lambda}_{i}(t,x)}{\lambda}\right),\quad j\in\mathbb{I}_{m}\setminus\{i\},\penalty 10000\ (t,x)\in\overline{{\cal D}}.$ (3.5)
Plugging (3.5) into (3.4), we get
$\frac{\partial V^{\lambda}_{i}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}V^{\lambda}_{i}(t,x)+f(t,x,i)+\lambda\sum_{j\neq i}\exp\left(\frac{V^{\lambda}_{j}(t,x)-g_{ij}-V^{\lambda}_{i}(t,x)}{\lambda}\right)=0,\penalty 10000\ (t,x)\in{\cal D},$ (3.6)
with the terminal condition $V^{\lambda}_{i}(T,x)=h(x)$ for $x\in\mathds{R}^{n}$.
To establish the well-posedness of the HJB system (3.4), we impose the following assumption.
Assumption 3.1.
The running reward function $f(\cdot,\cdot,i)\in C^{\alpha}({\cal D})$ for $i\in\mathbb{I}_{m}$ and terminal reward function $h(\cdot)\in C^{2+\alpha}({\cal D})$.
Lemma 3.2.
Let Assumptions 2.1, 2.2 and 3.1 hold. Then for any $\lambda>0$, the system of HJB equations (3.4) has a classical solution $(V^{\lambda}_{1},V^{\lambda}_{2},\cdots,V^{\lambda}_{m})$ with $V^{\lambda}_{i}\in C^{1,2}({\cal D})\cap C(\overline{{\cal D}})$ for $i\in\mathbb{I}_{m}$.
Proof.
Given $M>0$, consider a smooth and non-decreasing cut-off function $\phi_{M}$ such that $\phi_{M}(x)=e^{x}$ for $x\leq M$, $\phi_{M}(x)\leq e^{M+1}$ for $x\in(M,M+1)$ and $\phi_{M}(x)=e^{M+1}$for $x\geq M+1$. Hence, $\phi_{M}$ is bounded and Lipschitz continuous. Denote ${\cal D}_{N}:=\{(t,x):(t,x)\in{\cal D},|x|<N\}$. First, we will solve the following PDE system: for $i\in\mathbb{I}_{m}$,
$\begin{cases}\displaystyle\frac{\partial V_{i}^{M,N}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}V_{i}^{M,N}(t,x)+f(t,x,i)+\lambda\sum_{j\neq i}\phi_{M}\left(\frac{V^{M,N}_{j}(t,x)-V^{M,N}_{i}(t,x)-g_{ij}}{\lambda}\right)=0,\\[10.00002pt] \displaystyle\hfill(t,x)\in{\cal D}_{N},\\[5.0pt] \displaystyle V_{i}^{M,N}(t,x)=K(T-t)+h(x),\quad(t,x)\in\partial{\cal D}_{N},\end{cases}$ (3.7)
where the constant $K>0$ is given by
$K:=K_{f,h}+\lambda\sup_{i\in\mathbb{I}_{m}}\left(\sum_{j\neq i}\exp\left(-\frac{g_{ij}}{\lambda}\right)\right).$ (3.8)
For $i\in\mathbb{I}_{m}$, let us introduce the function
$u_{i}(t,x)$ $=K(T-t)+K_{f,h},\quad(t,x)\in\overline{{\cal D}}_{N}.$
It follows from Assumption 2.2 that
$\frac{\partial u_{i}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}u_{i}(t,x)+f(t,x,i)+\lambda\sum_{j\neq i}\phi_{M}\left(\frac{u_{j}(t,x)-u_{i}(t,x)-g_{ij}}{\lambda}\right)$
$=-K+f(t,x,i)+\lambda\sum_{j\neq i}\phi_{M}\left(-\frac{g_{ij}}{\lambda}\right)\leq 0,\quad\forall(t,x)\in{\cal D}_{N},$ (3.9)
and $u_{i}(t,x)\geq V^{\lambda}_{i}(t,x)$ for all $(t,x)\in\partial{\cal D}_{N}$. Similarly, we have
$\frac{\partial(-u_{i}(t,x))}{\partial t}+\mathcal{L}^{i}_{x}(-u_{i}(t,x))+f(t,x,i)+\lambda\sum_{j\neq i}\phi_{M}\left(\frac{(-u_{j}(t,x))-(-u_{i}(t,x))-g_{ij}}{\lambda}\right)$
$=K+f(t,x,i)+\lambda\sum_{j\neq i}\phi_{M}\left(-\frac{g_{ij}}{\lambda}\right)\geq 0,\quad\forall(t,x)\in{\cal D}_{N},$ (3.10)
and $u_{i}(t,x)\leq V_{i}^{\lambda}(t,x)$ for all $(t,x)\in\partial{\cal D}_{N}$. Invoking Theorem 2.1 in kusano1965first, we obtain that system (3.7) has a classical solution $(V_{1}^{M,N},\cdots,V_{m}^{M,N})$, with $V_{i}^{M,N}\in C^{1+\delta}(\overline{{\cal D}}_{N})$ for any $\delta\in(0,1)$ and $V_{i}^{M,N}\in C^{2+\alpha}(\overline{{\cal D}}_{N})$. Furthermore, we deduce from the comparison theorem (Theorem 1.3 in kusano1965first) that
$|V_{i}^{M,N}(t,x)|\leq u_{i}(t,x)=K(T-t)+K_{f,h},\quad\forall(i,t,x)\in\mathbb{I}_{m}\times\overline{{\cal D}}_{N},$ (3.11)
which implies that $V_{i}^{M,N}(t,x)$ is bounded. Thus, by choosing some $M$ large enough, for each $i\in\mathbb{I}_{m}$, $V^{N}_{i}:=V^{M,N}_{i}$ solves the following PDE
$\begin{cases}\displaystyle\frac{\partial V_{i}^{N}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}V_{i}^{N}(t,x)+f(t,x,i)+\lambda\sum_{j\neq i}\exp\left(\frac{V^{N}_{j}(t,x)-V^{N}_{i}(t,x)-g_{ij}}{\lambda}\right)=0,\quad(t,x)\in{\cal D}_{N},\\[10.00002pt] \displaystyle V_{i}^{N}(t,x)=K(T-t)+h(x),\quad(t,x)\in\partial{\cal D}_{N}.\end{cases}$ (3.12)
First, we apply Lemma 2 in kusano1965first to the problem (3.12) to derive for any $\delta\in(0,1)$,
[…]
In particular, $V_{i}^{N}||_{C^{\alpha}({\cal D}_{N})}$ are bounded independently of $N$. We then apply Lemma 1 in kusano1965first to the problem (3.12), obtaining
[…]
Consequently, we can extract from $\{V_{i}^{N}(t,x)\}$ a subsequence converging uniformly in $\overline{{\cal D}}$ together with the first $x$ , $t$-derivatives and second $x$-derivatives to a limit function $V_{i}^{\lambda}$, which is a solution to the HJB system (3.4). Thus, we complete the proof of the theorem. ∎
By the proof of Lemma 3.2, for any $\lambda>0$, the system of HJB equations (3.4) admits a classical solution $(V^{\lambda}_{1},V^{\lambda}_{2},\cdots,V^{\lambda}_{m})$ satisfying
$|V_{i}^{\lambda}(t,x)|\leq K(T-t)+K_{f,h},\quad\forall(i,t,x)\in\mathbb{I}_{m}\times\overline{{\cal D}},$ (3.13)
where the constant $K>0$ is given by (3.8). We now prove that this bounded classical solution is unique and coincides with the value function.
Proposition 3.3 (Verification Theorem).
Suppose Assumptions 2.1, 2.2, and 3.1 hold, and let $(V_{1}^{\lambda},V_{2}^{\lambda},\cdots,V_{m}^{\lambda})$ be a bounded classical solution to system (3.4), as provided by Lemma 3.2. We define a set of feedback functions by
$\pi_{ij}^{*}(t,x)=\exp\left(\frac{V^{\lambda}_{j}(t,x)-g_{ij}-V^{\lambda}_{i}(t,x)}{\lambda}\right),\quad j\in\mathbb{I}_{m}\setminus\{i\},\penalty 10000\ (t,x)\in\overline{{\cal D}},$ (3.14)
and
$\pi_{ii}^{*}(t,x)=-\sum_{j\neq i}\pi_{ij}^{*}(t,x),\quad(t,x)\in\overline{{\cal D}}.$ (3.15)
Consider the process $X^{*}$ governed by the dynamics (3.1), where the generator of the process $I^{*}$ is given by ${\bm{\pi}}^{*}=(\pi^{ij,*}_{t})_{i,j\in\mathbb{I}_{m},t\in[0,T]}$ and $\pi_{t}^{ij,*}=\pi_{ij}^{*}(t,X_{t}^{*})$. Then, for each $i\in\mathbb{I}_{m}$, the function $V_{i}^{\lambda}$ is the value function for problem (3.3), and the policy $\bm{\pi}$$^{*}$ is optimal.
Proof.
For $(i,t,x)\in\mathbb{I}_{m}\times\overline{{\cal D}}$, ${\bm{\pi}}\in\mathbb{U}_{t}$ and $s\in[t,T]$, using Itô’s rule, we obtain
$V_{I_{s}}^{\lambda}(s,X_{s})$ $=V_{i}^{\lambda}(t,x)+\int_{t}^{s}\left(\frac{\partial V_{I_{\ell}}^{\lambda}(\ell,X_{\ell})}{\partial t}+\mathcal{L}^{i}_{x}V_{I_{\ell}}^{\lambda}(\ell,X_{\ell})\right)d\ell+\int_{t}^{s}(D_{x}V_{I_{\ell}}^{\lambda}(\ell,X_{\ell}))^{\top}\sigma(\ell,X_{\ell},I_{\ell})dW_{\ell}$
$\quad+\int_{t}^{s}\sum_{j\neq I_{\ell}}\left(\pi_{\ell}^{ij}(V^{\lambda}_{j}(\ell,X_{\ell})-V^{\lambda}_{I_{\ell}}(\ell,X_{\ell}))\right)d\ell.$ (3.16)
Taking the expectation on both sides of Eq. (3), it follows from (3.4) that
$V_{i}^{\lambda}(t,x)\geq\mathbb{E}_{t,x,i}\bigg[\int^{s}_{t}f(\ell,X_{\ell},I_{\ell})ds-\sum_{k=1}^{\infty}g_{\kappa_{k-1}\kappa_{k}}{\bf 1}_{\{\tau_{k}\leq s\}}+\lambda\int^{s}_{t}R({\bm{\pi}}_{\ell},I_{\ell})ds+V_{I_{s}}^{\lambda}(s,X_{s})\bigg].$ (3.17)
Letting $s\to T$ in (3.17), we get from $V_{i}^{\lambda}=h(i,t,x)$ and the dominated convergence theorem that
$V_{i}^{\lambda}(t,x)\geq\mathbb{E}_{t,x,i}\bigg[\int^{T}_{t}f(\ell,X_{\ell},I_{\ell})ds-\sum_{k=1}^{\infty}g_{\kappa_{k-1}\kappa_{k}}{\bf 1}_{\{\tau_{k}\leq T\}}+\lambda\int^{T}_{t}R({\bm{\pi}}_{\ell},I_{\ell})ds+h(X_{T})\bigg].$ (3.18)
The inequality (3.18) holds for any ${\bm{\pi}}\in\mathbb{U}_{t}$ and becomes an equality when ${\bm{\pi}}={\bm{\pi}^{*}}$. Furthermore, Theorem 2.6 in nguyen2025hybrid guarantees the existence and uniqueness of the strong solution $(X^{*},I^{*})$ to the SDE (3.1). Thus, we complete the proof of the theorem. ∎
4 Policy Iteration and Convergence
The goal of this section is to study the policy iteration using the characterization in (3.5). In particular, in the context of optimal regime switching, we aim to show the policy improvement and the convergence of policy iterations, which demonstrate that each policy update guarantees the performance enhancement and the repeated iterations will lead to the desired optimal policy when the model is known. We also examine the connection between our exploratory formulation and the classical optimal switching problem by analyzing the limit of the vanishing regularization.
We first focus on the rule of policy iteration. Given a feedback strategy ${\bm{\pi}}^{n}(t,x)=(\pi_{ij}^{n}(t,x))_{i,j\in\mathbb{I}_{m}}$, the corresponding value function $(V^{n}_{1},\cdots,V^{n}_{m})$ satisfies the following PDE system: for $i\in\mathbb{I}_{m}$,
$\begin{cases}\displaystyle\frac{\partial V_{i}^{n}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}V_{i}^{n}(t,x)+f(t,x,i)+H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n}(t,x),\cdots,V_{m}^{n}(t,x))=0,\\ \displaystyle V_{i}^{n}(T,x)=h(x).\end{cases}$ (4.1)
Here, the Hamiltomian $H_{i}({\bm{\pi}}_{i},{\bm{y}}):\mathds{R}^{m}\times\mathds{R}^{m}\to\mathds{R}$ is defined by
$H_{i}({\bm{\pi}}_{i},{\bm{y}})=\sum_{j\neq i}\pi_{ij}(y_{j}-g_{ij}-y_{i})+\lambda\sum_{j\neq i}(\pi_{ij}-\pi_{ij}\log\pi_{ij}).$ (4.2)
Having the value function pair $(V^{n}_{1},\cdots,V_{m}^{n})$, one can construct a feedback strategy ${\bm{\pi}}^{n+1}$ satisfying
$\pi_{ij}^{n+1}(t,x)=\exp\left(\frac{V^{n}_{j}(t,x)-g_{ij}-V^{n}_{i}(t,x)}{\lambda}\right),\penalty 10000\ i,j\in\mathbb{I}_{m},j\neq i.$ (4.3)
We continue this iteration, generating a sequence of strategy-value function pairs. The following theorem states that each iteration improves the value function.
Proposition 4.1.
Let Assumptions 2.1, 2.2 and 3.1 hold. Give any initial guess $(V^{0}_{1},\cdots,V^{0}_{m})$ with $V^{0}_{i}\in C^{0}(\overline{{\cal D}})$ for $i\in\mathbb{I}_{m}$. $\{(V^{n}_{i},\pi^{n}_{ij})_{i,j\in\mathbb{I}_{m}}\}_{n=1,2,\ldots}$ are defined iteratively according to (4.1) and (4.3). Then, we have that $V^{n}_{i}\leq V^{n+1}_{i}\leq V_{i}^{\lambda}$ for $i\in\mathbb{I}_{m}$ and $n=1,2,\ldots$.
Proof.
By Proposition 3.3, the tuple $(V^{\lambda}_{1},\cdots,V_{m}^{\lambda})$ constitutes the optimal value function for the control problem (3.2). Since $(V^{n}_{1},\cdots,V^{n}_{m})$ is the value function corresponding to the strategy ${\bm{\pi}}^{n}(t,x)$ for the same problem, it follows that $V^{n}_{i}\leq V_{i}^{\lambda}$ for all $i\in\mathbb{I}_{m}$ and $n\geq 1$.
Given the uniform bound $V^{n}_{i}\leq V_{i}^{\lambda}$, the estimate (3.13), $V^{0}_{i}\in C^{0}(\overline{{\cal D}})$, and the policy iteration definition (4.3), we deduce that for any $n\geq 1$ and $i,j\in\mathbb{I}_{m}$, the function $\pi_{ij}^{n}(t,x)$ is bounded on $\overline{{\cal D}}$. For $n\geq 1$, proceeding with a truncation argument analogous to the proof of Lemma 3.1, we select a suitable constant $\tilde{K}$ and consider the following pair of PDE systems defined on the truncated domain ${\cal D}_{N}$:
$\begin{cases}\displaystyle\frac{\partial V_{i}^{n,N}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}V_{i}^{n,N}(t,x)+f(t,x,i)\\ \displaystyle\quad+\,H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n,N}(t,x),\cdots,V_{m}^{n,N}(t,x))=0,&(t,x)\in{\cal D}_{N},\\[10.00002pt] \displaystyle V_{i}^{n,N}(t,x)=\tilde{K}(T-t)+h(x),&(t,x)\in\partial{\cal D}_{N},\end{cases}$ (4.4)
and
$\begin{cases}\displaystyle\frac{\partial V_{i}^{n+1,N}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}V_{i}^{n+1,N}(t,x)+f(t,x,i)\\ \displaystyle\quad+\,H_{i}({\bm{\pi}}^{n+1,N}_{i}(t,x),V_{1}^{n+1,N}(t,x),\cdots,V_{m}^{n+1,N}(t,x))=0,&(t,x)\in{\cal D}_{N},\\[10.00002pt] \displaystyle V_{i}^{n+1,N}(t,x)=\tilde{K}(T-t)+h(x),&(t,x)\in\partial{\cal D}_{N},\end{cases}$ (4.5)
where policy ${\bm{\pi}}^{n+1,N}$ is given by
$\pi_{ij}^{n+1,N}(t,x)=\exp\left(\frac{V^{n,N}_{j}(t,x)-g_{ij}-V^{n,N}_{i}(t,x)}{\lambda}\right),\penalty 10000\ i,j\in\mathbb{I}_{m},\penalty 10000\ j\neq i.$ (4.6)
As in Lemma 3.1, the bounded solutions satisfy $V^{n,N}_{i}\to V^{n}_{i}$ and $V^{n+1,N}_{i}\to V^{n+1}_{i}$ as $N\to\infty$ for each $i\in\mathbb{I}_{m}$.
Define the difference $\Delta_{i}^{n,N}(t,x):=V_{i}^{n+1,N}(t,x)-V_{i}^{n,N}(t,x)$ for $i\in\mathbb{I}_{m}$ and $(t,x)\in\overline{{\cal D}}_{N}$. Subtracting (4.4) from (4.5) yields the equation satisfied by $\Delta_{i}^{n,N}$:
$\frac{\partial\Delta_{i}^{n,N}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}\Delta_{i}^{n,N}(t,x)$ $+H_{i}({\bm{\pi}}^{n+1,N}_{i}(t,x),V_{1}^{n+1,N}(t,x),\cdots,V_{m}^{n+1,N}(t,x))$
$-H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n,N}(t,x),\cdots,V_{m}^{n,N}(t,x))=0,\quad(t,x)\in{\cal D}_{N},$ (4.7)
with boundary condition $\Delta_{i}^{n,N}(t,x)=0$ for $(t,x)\in\partial{\cal D}_{N}$. A key observation from (4.6) is that ${\bm{\pi}}_{i}^{n+1,N}(t,x)$ maximizes the Hamiltonian:
${\bm{\pi}}_{i}^{n+1,N}(t,x)=\underset{{\bm{\pi}}}{\mathrm{arg\,max\,}}H_{i}({\bm{\pi}},V_{1}^{n,N}(t,x),\cdots,V_{m}^{n,N}(t,x)).$ (4.8)
Using (4) and the maximizing property (4.8), we derive for $(t,x)\in{\cal D}_{N}$:
$\frac{\partial\Delta_{i}^{n,N}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}\Delta_{i}^{n,N}(t,x)+\sum_{j\neq i}\pi_{ij}^{n+1,N}(t,x)\Delta_{j}^{n,N}(t,x)-\sum_{j\neq i}\pi_{ij}^{n+1,N}(t,x)\Delta_{i}^{n,N}(t,x)$
$=-H_{i}({\bm{\pi}}_{i}^{n+1,N}(t,x),V_{1}^{n+1,N}(t,x),\cdots,V_{m}^{n+1,N}(t,x))+\sum_{j\neq i}\pi_{ij}^{n+1,N}(t,x)(\Delta_{j}^{n,N}(t,x)-\Delta_{i}^{n,N}(t,x))$
$\quad+H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n,N}(t,x),\cdots,V_{m}^{n,N}(t,x))$
$=H_{i}({\bm{\pi}}^{n}_{i}(t,x),V_{1}^{n,N}(t,x),\cdots,V_{m}^{n,N}(t,x))-H_{i}({\bm{\pi}}^{n+1,N}_{i}(t,x),V_{1}^{n,N}(t,x),\cdots,V_{m}^{n,N}(t,x))$
$\leq 0.$ (4.9)
Applying Theorem 1.3 in kusano1965first, we deduce that $\Delta^{n,N}_{i}(t,x)\geq 0$, that is, $V_{i}^{n+1,N}(t,x)\geq V_{i}^{n,N}(t,x)$, for all $i\in\mathbb{I}_{m}$ and $(t,x)\in\overline{{\cal D}}$. Passing to the limit as $N\to\infty$, we have $V_{i}^{n+1}(t,x)\geq V_{i}^{n}(t,x)$, which then completes the proof. ∎
The following theorem, as the first main result of this paper, establishes a fundamental convergence guarantee for our policy iteration method, demonstrating that the sequence of value functions $(V_{1}^{n},\cdots,V_{m}^{n})$ generated through successive iterations converges uniformly to the optimal value functions $(V_{1}^{\lambda}\cdots,V^{\lambda}_{m})$ of our exploratory optimal switching problem. Moreover, we can obtain the explicit convergence rate for the policy iteration.
Theorem 4.2.
Let Assumptions 2.1, 2.2 and 3.1 hold. Give any initial guess $(V^{0}_{1},\cdots,V^{0}_{m})$ with $V^{0}_{i}\in C^{0}(\overline{{\cal D}})$ for $i\in\mathbb{I}_{m}$. $\{(V^{n}_{i},\pi^{n}_{ij})_{i,j\in\mathbb{I}_{m}}\}_{n=1,2,\ldots}$ are defined iteratively according to (4.1) and (4.3). Then, we have that, for all $n\geq 1$,
$\sup_{i\in\mathbb{I}_{m}}\sup_{(t,x)\in\overline{{\cal D}}}|V_{i}^{n}(t,x)-V_{i}^{\lambda}(t,x)|\leq C_{1}\frac{C_{2}^{n}}{n!},$ (4.10)
where $C_{1},C_{2}>0$ are constants independent of $n$.
Proof.
For $n\geq 0$, let us introduce the function $F^{n}:[0,T]\to\mathds{R}_{+}$ given by
$F^{n}(t):=\sup_{i\in\mathbb{I}_{m}}\sup_{x\in\mathds{R}^{n}}|V_{i}^{n}(t,x)-V_{i}^{\lambda}(t,x)|.$ (4.11)
By the proof of Lemma 3.2, we can obtain
$|V_{i}^{\lambda}(t,x)|\leq K(T-t)+K_{f,h}\leq KT+K_{f,h},\quad\forall(i,t,x)\in\mathbb{I}_{m}\times\overline{{\cal D}},$ (4.12)
where the constant $K$ is given by (3.8). This implies the boundedness of $V_{i}^{\lambda}(t,x)$, which in turn implies that the policy $\bm{\pi}$$^{*}$ from (3.5) is bounded. Similarly, by using Theorem 4.1 and (4.3), we can deduce that the sequence of functions $V^{n}_{i}(t,x)$ and the corresponding policies ${\bm{\pi}}^{n}(t,x)$ are uniformly bounded for $n\geq 1$. Then, it follows from (3.5), (4.2) and (4.3) that
$\left|H_{i}({\bm{\pi}}_{i}^{n},V_{1}^{\lambda},\cdots,V_{m}^{\lambda})-H_{i}({\bm{\pi}}_{i}^{*},V_{1}^{\lambda},\cdots,V_{m}^{\lambda})\right|$
$=\Bigg|\sum_{j\neq i}\pi_{ij}^{n}(V_{j}^{\lambda}-g_{ij}-V_{i}^{\lambda})+\lambda\sum_{j\neq i}(\pi_{ij}^{n}-\pi_{ij}^{n}\log\pi_{ij}^{n})$
$\quad+\left(\sum_{j\neq i}\pi_{ij}^{*}(V_{j}^{\lambda}-g_{ij}-V_{i}^{\lambda})+\lambda\sum_{j\neq i}(\pi_{ij}^{*}-\pi_{ij}^{*}\log\pi_{ij}^{*})\right)\Bigg|$
$\leq\sum_{j\neq i}\pi^{n}_{ij}|V_{j}^{\lambda}-V_{j}^{n}|+|V_{i}^{\lambda}-V_{i}^{n}|\sum_{j\neq i}\pi^{n}_{ij}$
$\quad+\lambda\sum_{j\neq i}\left|\exp\left(\frac{V^{n}_{j}-g_{ij}-V^{n}_{i}}{\lambda}\right)-\exp\left(\frac{V_{j}^{\lambda}-g_{ij}-V_{i}^{\lambda}}{\lambda}\right)\right|$
$\leq C^{*}F^{n}(t),$ (4.13)
where $C^{*}>0$ is a constant independent of $n$. For $n\geq 0$, we define the function $w_{i}^{n}:\overline{{\cal D}}\to\mathds{R}$ for $i\in\mathbb{I}_{m}$ as
$w_{i}^{n}(t,x):=V_{i}^{\lambda}(t,x)-V_{i}^{n+1}(t,x)-C^{*}\int_{t}^{T}F^{n}(s)ds,\quad(t,x)\in\overline{{\cal D}}.$
By using (4), it holds that for any $(t,x)\in{\cal D}$,
$\frac{\partial w_{i}^{n}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}w_{i}^{n}(t,x)+\sum_{j\neq i}\pi_{ij}^{n+1}(t,x)w_{j}^{n}(t,x)-\sum_{j\neq i}\pi_{ij}^{n+1}(t,x)w_{i}^{n}(t,x)$
$=H_{i}({\bm{\pi}}_{i}^{n+1}(t,x),V_{1}^{\lambda}(t,x),\cdots,V_{m}^{\lambda}(t,x))-H_{i}({\bm{\pi}}^{*}_{i}(t,x),V_{1}^{\lambda}(t,x),\cdots,V_{m}^{\lambda}(t,x))+C^{*}F^{n}(t)\geq 0,$
and $w_{i}^{n}(T,x)=0$ for $x\in\mathds{R}^{n}$. By virtue of Theorem 1.3 in kusano1965first, we deduce $w^{n}_{i}(t,x)\leq 0$. That is,
$V_{i}^{\lambda}(t,x)-V_{i}^{n+1}(t,x)\leq C^{*}\int_{t}^{T}F^{n}(s)ds,\quad\forall(i,t,x)\in\mathbb{I}_{m}\times\overline{{\cal D}}.$ (4.14)
This yields the inequality
$F^{n+1}(t)\leq C^{*}\int_{t}^{T}F^{n}(s)ds,\quad\forall t\in[0,T],$ (4.15)
from which we deduce that
$F^{n}(t)\leq\frac{(C^{*})^{n}T^{n}}{n!}\sup_{s\in[0,T]}F^{1}(s),\quad\forall t\in[0,T].$ (4.16)
Because $F^{1}(t)$ is bounded, let $C_{2}=C^{*}T$ and $C_{1}=\sup_{t\in[0,T]}F^{1}(t)$. Then we obtain that desired result. ∎
To establish a connection between our exploratory formulation and the classical optimal switching problem, we next rigorously analyze the convergence result of the exploratory solution as the temperature parameter $\lambda$ approaches zero. Unlike the existing results in tang2022exploratory for regular control problem that focus on a single PDE problem, the nature of problem with multiple regime states calls for some distinct analysis to investigate the system of PDEs in our setting. In particular, we employ some stability analysis of viscosity solutions to the PDE system to examine the limit of vanishing entropy regularization. The mathematical goal is to show that the solution of the system of PDE will converge to the solution of the system of variational inequalities as $\lambda\rightarrow 0$.
Let us introduce the upper and lower weak limits of functions $(V_{1}^{\lambda},\cdots,V_{m}^{\lambda})$ defined as follows: for $i\in\mathbb{I}_{m}$ and $(t,x)\in\overline{{\cal D}}$,
$\overline{V}_{i}(t,x):=\begin{cases}\displaystyle\limsup_{\lambda\to 0\atop(s,y)\to(t,x),(s,y)\in{\cal D}}V_{i}^{\lambda}(s,y),&(t,x)\in{\cal D},\\[10.00002pt] \displaystyle\qquad h(x),&t=T,\penalty 10000\ x\in\mathds{R}^{n},\end{cases}$ (4.17)
and
$\underline{V}_{i}(t,x):=\begin{cases}\displaystyle\liminf_{\lambda\to 0,\atop(s,y)\to(t,x),(s,y)\in{\cal D}}V_{i}^{\lambda}(s,y),&(t,x)\in{\cal D},\\[10.00002pt] \displaystyle\qquad h(x),&t=T,\penalty 10000\ x\in\mathds{R}^{n}.\end{cases}$ (4.18)
The next lemma plays a crucial role in establishing the convergence of the value functions $(V_{1}^{\lambda},\cdots,V_{m}^{\lambda})$ as the temperature parameter $\lambda$ tends to zero. By defining the upper and lower weak limits, we capture the limiting behavior of these functions. The result asserts that these limits are bounded and satisfy the viscosity solution properties for the system of HJB equations (2.7). Specifically, the upper weak limits form a viscosity subsolution, and the lower weak limits form a viscosity supersolution.
Lemma 4.3.
Let Assumptions 2.1, 2.2, and 3.1 hold. Consider the upper and lower weak limits of the functions $(V_{1}^{\lambda},\cdots,V_{m}^{\lambda})$, defined by (4.17) and (4.18), respectively. Then the tuple of upper weak limits $(\overline{V}_{1},\cdots,\overline{V}_{m})$ is a bounded viscosity subsolution of system (2.7), while the tuple of lower weak limits $(\underline{V}_{1},\cdots,\underline{V}_{m})$ is a bounded viscosity supersolution of system (2.7).
Proof.
It follows from (3.8), (3.13) and Assumption 2.2-(ii) that
$|V_{i}^{\lambda}(t,x)|\leq KT+K_{f,h}$ $=K_{f,h}T+\lambda\sup_{i\in\mathbb{I}_{m}}\left(\sum_{j\neq i}\exp\left(-\frac{g_{ij}}{\lambda}\right)\right)T+K_{f,h}$
$\leq K_{f,h}T+\lambda(m-1)T+K_{f,h}$
for all $\lambda>0$ and $(i,t,x)\in\mathbb{I}_{m}\times\overline{{\cal D}}$. This implies that $\overline{V}_{i}$ and $\underline{V}_{i}$ for $i\in\mathbb{I}_{m}$ are bounded functions. Applying Lemma 1.5 in Chapter V of bardi1997optimal, $\overline{V}_{i}$ is upper-semicontinuous on $\cal D$} while $\underline{V}_{i}$ is lower-semicontinuous on $\cal D$} for every $i\in\mathbb{I}_{m}$.
We next show that the tuple of upper weak limits $(\overline{V}_{1},\cdots,\overline{V}_{m})$ is a viscosity subsolution of system (2.7) using the contradiction argument. For $i\in\mathbb{I}_{m}$, let $(t_{0},x_{0})\in{\cal D}$ and the test function $\varphi_{i}\in C^{1,2}({\cal D})$ such that $(t_{0},x_{0})$ is a local maximum of $\overline{V}_{i}-\varphi_{i}$. Assume that
$\min\Bigg\{$ $-\frac{\partial\varphi_{i}(t_{0},x_{0})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{0},x_{0})-f(t_{0},x_{0},i),$
$\qquad\qquad\quad\overline{V}_{i}(t_{0},x_{0})-\max_{j\neq i}(\overline{V}_{j}(t_{0},x_{0})-g_{ij})\Bigg\}>0.$ (4.19)
That is,
$\delta:=-\frac{\partial\varphi_{i}(t_{0},x_{0})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{0},x_{0})-f(t_{0},x_{0},i)>0,$ (4.20)
$\varepsilon:=\overline{V}_{i}(t_{0},x_{0})-\max_{j\neq i}(\overline{V}_{j}(t_{0},x_{0})-g_{ij})>0.$ (4.21)
In view of Lemma 1.6 in Chapter V of bardi1997optimal, there exists a sequence $\{(t_{n},x_{n})\}_{n\geq 1}$ with $(t_{n},x_{n})\in{\cal D}$ and a sequence $\{\lambda_{n}\}_{n\geq 1}$ with $\lambda_{n}>0$, $\lim_{n\to\infty}\lambda_{n}=0$ such that $(t_{n},x_{n})$ is a local maximum point of $V^{\lambda_{n}}_{i}-\varphi_{i}$ and
$\lim_{n\to\infty}(t_{n},x_{n})=(t_{0},x_{0}),\quad\lim_{n\to\infty}V^{\lambda_{n}}_{i}(t_{n},x_{n})=\overline{V}_{i}(t_{0},x_{0}).$ (4.22)
Lemma 3.2 implies that for any $\lambda>0$, $(V^{\lambda}_{1},V^{\lambda}_{2},\cdots,V^{\lambda}_{m})$ is a classical solution to the system of of HJB equations (3.4), thus $V^{\lambda_{n}}_{i}$ is a viscosity subsolution of the following PDE:
$-\frac{\partial V^{\lambda_{n}}_{i}(t,x)}{\partial t}-\mathcal{L}^{i}_{x}V^{\lambda_{n}}_{i}(t,x)-f(t,x,i)-\lambda\sum_{j\neq i}\exp\left(\frac{V^{\lambda_{n}}_{j}(t,x)-g_{ij}-V^{\lambda_{n}}_{i}(t,x)}{\lambda}\right)=0.$
Consequently, we have
$-\frac{\partial\varphi_{i}(t_{n},x_{n})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{n},x_{n})-f(t_{n},x_{n},i)-\lambda_{n}\sum_{j\neq i}\exp\left(\frac{V^{\lambda_{n}}_{j}(t_{n},x_{n})-g_{ij}-V^{\lambda_{n}}_{i}(t_{n},x_{n})}{\lambda}\right)\leq 0$ (4.23)
for any $n\geq 1$.
From (4.20), (4.21) and (4.22), it follows that there exists some $n_{1}>0$ such that for all $n\geq n_{1}$,
$-\frac{\partial\varphi_{i}(t_{n},x_{n})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{n},x_{n})-f(t_{n},x_{n},i)\geq\frac{\delta}{2},$
and for any $j\in\mathbb{I}_{m},j\neq i$,
$V^{\lambda_{n}}_{j}(t_{n},x_{n})-g_{ij}-V^{\lambda_{n}}_{i}(t_{n},x_{n})\leq-\frac{\varepsilon}{2}.$
Selecting $n_{2}$ such that for all $n\geq n_{2}$, $\lambda_{n}\exp(-\frac{\varepsilon}{2\lambda_{n}})<\frac{\delta}{2(m-1)}$, then for $n\geq\max\{n_{1},n_{2}\}$, we get that
$-\frac{\partial\varphi_{i}(t_{n},x_{n})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{n},x_{n})-f(t_{n},x_{n},i)-\lambda_{n}\sum_{j\neq i}\exp\left(\frac{V^{\lambda_{n}}_{j}(t_{n},x_{n})-g_{ij}-V^{\lambda_{n}}_{i}(t_{n},x_{n})}{\lambda}\right)$
$\geq-\frac{\partial\varphi_{i}(t_{n},x_{n})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{n},x_{n})-f(t_{n},x_{n},i)-\lambda_{n}\sum_{j\neq i}\exp\left(-\frac{\varepsilon}{2\lambda_{n}}\right)$
$\geq\frac{\delta}{2}-\lambda_{n}(m-1)\exp\left(-\frac{\varepsilon}{2\lambda_{n}}\right)>0.$ (4.24)
The inequalities (4.23) and (4) are contradictory. Therefore, we conclude that the assumption (4) is note true, which implies that $(\overline{V}_{1},\cdots,\overline{V}_{m})$ is a viscosity subsolution of system (2.7).
We next show that the tuple of lower weak limits $(\underline{V}_{1},\cdots,\underline{V}_{m})$ is a viscosity supersolution of system (2.7) by contradiction. For $i\in\mathbb{I}_{m}$, let $(t_{0},x_{0})\in{\cal D}$ and the test function $\varphi_{i}\in C^{1,2}({\cal D})$ such that $(t_{0},x_{0})$ is a local minimum of $\overline{V}_{i}-\varphi_{i}$. Assume that
$\min\Bigg\{$ $-\frac{\partial\varphi_{i}(t_{0},x_{0})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{0},x_{0})-f(t_{0},x_{0},i),$
$\qquad\qquad\quad\underline{V}_{i}(t_{0},x_{0})-\max_{j\neq i}(\underline{V}_{j}(t_{0},x_{0})-g_{ij})\Bigg\}<0.$ (4.25)
Using Lemma 1.6 in Chapter V of bardi1997optimal again, there exists a sequence $\{(t_{n},x_{n})\}_{n\geq 1}$ with $(t_{n},x_{n})\in{\cal D}$ and a sequence $\{\lambda_{n}\}_{n\geq 1}$ with $\lambda_{n}>0$, $\lim_{n\to\infty}\lambda_{n}=0$ such that $(t_{n},x_{n})$ is a local minimum point of $V^{\lambda_{n}}_{i}-\varphi_{i}$ and
$\lim_{n\to\infty}(t_{n},x_{n})=(t_{0},x_{0}),\quad\lim_{n\to\infty}V^{\lambda_{n}}_{i}(t_{n},x_{n})=\underline{V}_{i}(t_{0},x_{0}).$ (4.26)
By Lemma 3.2, for any $\lambda>0$, $(V^{\lambda}_{1},V^{\lambda}_{2},\cdots,V^{\lambda}_{m})$ is a classical solution to the system of of HJB equations (3.4), thus $V^{\lambda_{n}}_{i}$ is a viscosity supersolution of the following PDE:
$-\frac{\partial V^{\lambda_{n}}_{i}(t,x)}{\partial t}-\mathcal{L}^{i}_{x}V^{\lambda_{n}}_{i}(t,x)-f(t,x,i)-\lambda\sum_{j\neq i}\exp\left(\frac{V^{\lambda_{n}}_{j}(t,x)-g_{ij}-V^{\lambda_{n}}_{i}(t,x)}{\lambda}\right)=0.$
Therefore we have
$-\frac{\partial\varphi_{i}(t_{n},x_{n})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{n},x_{n})-f(t_{n},x_{n},i)-\lambda_{n}\sum_{j\neq i}\exp\left(\frac{V^{\lambda_{n}}_{j}(t_{n},x_{n})-g_{ij}-V^{\lambda_{n}}_{i}(t_{n},x_{n})}{\lambda}\right)\geq 0$ (4.27)
for any $n\geq 1$. We consider two cases for the inequality (4).
Case 1. Assume that
$-\frac{\partial\varphi_{i}(t_{0},x_{0})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{0},x_{0})-f(t_{0},x_{0},i)<0.$ (4.28)
By (4.27), we have
$-\frac{\partial\varphi_{i}(t_{n},x_{n})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{n},x_{n})-f(t_{n},x_{n},i)\geq\lambda_{n}\sum_{j\neq i}\exp\left(\frac{V^{\lambda_{n}}_{j}(t_{n},x_{n})-g_{ij}-V^{\lambda_{n}}_{i}(t_{n},x_{n})}{\lambda}\right)\geq 0,$
which yields
$-\frac{\partial\varphi_{i}(t_{0},x_{0})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{0},x_{0})-f(t_{0},x_{0},i)$
$=\lim_{n\to\infty}\left(-\frac{\partial\varphi_{i}(t_{n},x_{n})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{n},x_{n})-f(t_{n},x_{n},i)\right)\geq 0.$
Thus, we obtain a contradiction.
Case 2. Assume that
$\delta:=-\frac{\partial\varphi_{i}(t_{0},x_{0})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{0},x_{0})-f(t_{0},x_{0},i)\geq 0,$ (4.29)
and
$\varepsilon:=-(\underline{V}_{i}(t_{0},x_{0})-\max_{j\neq i}(\underline{V}_{j}(t_{0},x_{0})-g_{ij}))=\underline{V}_{k}(t_{0},x_{0})-g_{ik}-\underline{V}_{i}(t_{0},x_{0})>0.$ (4.30)
By (4.26), (4.29) and (4.30), there exists some $n_{1}>0$ such that for all $n\geq n_{1}$,
$-\frac{\partial\varphi_{i}(t_{n},x_{n})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{n},x_{n})-f(t_{n},x_{n},i)\leq\frac{3\delta}{2},$
and
$V^{\lambda_{n}}_{k}(t_{n},x_{n})-g_{ik}-V^{\lambda_{n}}_{i}(t_{n},x_{n})\geq\frac{\varepsilon}{2}.$
Selecting $n_{2}$ such that for all $n\geq n_{2}$, $\lambda_{n}\exp(\frac{\varepsilon}{2\lambda_{n}})>\frac{3\delta}{2}$, then for $n\geq\max\{n_{1},n_{2}\}$, it holds that
$-\frac{\partial\varphi_{i}(t_{n},x_{n})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{n},x_{n})-f(t_{n},x_{n},i)-\lambda_{n}\sum_{j\neq i}\exp\left(\frac{V^{\lambda_{n}}_{j}(t_{n},x_{n})-g_{ij}-V^{\lambda_{n}}_{i}(t_{n},x_{n})}{\lambda}\right)$
$\leq-\frac{\partial\varphi_{i}(t_{n},x_{n})}{\partial t}-\mathcal{L}^{i}_{x}\varphi_{i}(t_{n},x_{n})-f(t_{n},x_{n},i)-\lambda_{n}\exp\left(\frac{V^{\lambda_{n}}_{k}(t_{n},x_{n})-g_{ik}-V^{\lambda_{n}}_{i}(t_{n},x_{n})}{\lambda}\right)$
$\leq\frac{3\delta}{2}-\lambda_{n}\exp\left(\frac{\varepsilon}{2\lambda_{n}}\right)<0.$ (4.31)
The inequalities (4.27) and (4) are contradictory.
Combining the arguments in two cases above, we conclude that assertion (4) does not hold. This implies that $(\overline{V}_{1},\cdots,\overline{V}_{m})$ is a viscosity supersolution of system (2.7), which completes the proof. ∎
As the second main result of this paper, the next theorem shows the convergence result towards the classical optimal switching problem as the entropy regularization vanishes.
Theorem 4.4.
Let Assumptions 2.1, 2.2 and 3.1 hold. Consider the value functions $(V_{1},\cdots,V_{m})$ of the classical optimal switching problem defined by (2.6), and the value functions $(V_{1}^{\lambda},\cdots,V_{m}^{\lambda})$ of the exploratory optimal switching problem defined by (3.3). Then for any $i\in\mathbb{I}_{m}$ and $(t,x)\in\overline{{\cal D}}$,
$\lim_{\lambda\to 0}V_{i}^{\lambda}(t,x)=V_{i}(t,x).$ (4.32)
Proof.
By using Lemma 4.3 and Lemma 2.3, we have
$\overline{V}_{i}(t,x)\leq\underline{V}_{i}(t,x),\quad\forall i\in\mathbb{I}_{m},(t,x)\in\overline{{\cal D}}.$
On the other hand, it follows from the definition of upper and lower weak limits that $\overline{V}_{i}(t,x)\geq\underline{V}_{i}(t,x)$, for any $i\in\mathbb{I}_{m}$ and $(t,x)\in\overline{{\cal D}}$. Thus, $\overline{V}_{i}(t,x)=\underline{V}_{i}(t,x)$, and we denote by
$V^{*}_{i}(t,x):=\overline{V}_{i}(t,x)=\underline{V}_{i}(t,x)\quad\text{for}\penalty 10000\ i\in\mathbb{I}_{m},(t,x)\in\overline{{\cal D}}.$
It follows from (4.17), (4.18) and Lemma 4.3 that $(V_{1}^{*},\cdots,V_{m}^{*})$ is a bounded viscosity solution of system (2.7) satisfying $V^{*}_{i}(t,x)=\lim_{\lambda\to 0}V_{i}^{\lambda}(t,x)$. We deduce from Theorem 2.4 that
$V_{i}(t,x)=V^{*}_{i}(t,x)=\lim_{\lambda\to 0}V_{i}^{\lambda}(t,x).$ (4.33)
Thus, we complete the proof of the theorem. ∎
Theorem 4.4 justifies the use of the exploratory formulation as a well-founded mathematical relaxation: as the exploration effect diminishes (as the temperature parameter $\lambda\to 0$), the value function of the exploratory formulation indeed converges towards the value function of the classical optimal switching problem. Mathematically speaking, it is interesting to observe that the solution to the system of PDEs will converge to the solution of system of variational inequalities. Therefore, our exploratory formulation can also be regarded as a penalization approach to study a system of variational inequalities, under which we only need to handle the existence and regularity of solution to a system of PDEs.
5 Reinforcement Learning Algorithm
In this section, we design a RL algorithm to solve the exploratory optimal switching problem when the model is unknown. The core of our approach lies in a key reformulation: we have transformed the original optimal switching problem into a standard optimal control problem where we control the generator of the finite-state Markov chain that characterizes the switching regimes. The primary distinction from classical problems is that the agent now actively controls the transition rates between regimes, adding a continuous layer of decision-making on top of the discrete switching choices.
Our choice of the randomization and the exploratory form leads to an explicit characterization of the optimal policy that depends on the value functions, without involving their derivatives. Leveraging this solution structure, we adopt the policy evaluation (PE) method based on the martingale characterization method similar to jia2022policy, which consider two alternative methods based on a martingale characterization: minimizing a martingale loss function, which provides the best mean-square approximation of the true value function, and solving a system of martingale orthogonality condition with test functions. In what follows, we design the PE algorithm by the martingale orthogonality condition and the established policy improvement result in Proposition 4.1.
Recall that given a feedback strategy ${\bm{\pi}}(t,x)=(\pi_{ij}(t,x))_{i,j\in\mathbb{I}_{m}}$, the corresponding value function $(v_{1}^{{\bm{\pi}}},\cdots,v_{m}^{{\bm{\pi}}})$ satisfies the PDE system that for $i\in\mathbb{I}_{m}$,
$\begin{cases}\displaystyle\frac{\partial v_{i}^{{\bm{\pi}}}(t,x)}{\partial t}+\mathcal{L}^{i}_{x}v_{i}^{{\bm{\pi}}}(t,x)+f(t,x,i)+H_{i}({\bm{\pi}}_{i}(t,x),v_{1}^{{\bm{\pi}}}(t,x),\cdots,v^{{\bm{\pi}}}_{m}(t,x))=0,\\ \displaystyle v^{{\bm{\pi}}}_{i}(T,x)=h(x),\end{cases}$ (5.1)
where the Hamiltomian $H_{i}$ is given by (4.2). For simplicity, we omit the superscript $\bm{\pi}$}} and denote the value function as:
$v(t,x,i)=v^{{\bm{\pi}}}_{i}(t,x),\quad\text{for }i\in\mathbb{I}_{m},\penalty 10000\ (t,x)\in\overline{{\cal D}},$ (5.2)
and denote by $I=(I_{t})_{t\geq 0}$ a continuous-time finite-state Markov chain with generator ${\bm{\pi}}=(\pi^{ij})_{i,j\in\mathbb{I}_{m}}$. Let us introduce the process $M=(M_{t})_{t\in[0,T]}$ given by
$M_{t}:=v(t,X_{t},I_{t})+\int_{0}^{t}\left(f(s,X_{s},I_{s})+\lambda R({\bm{\pi}}_{s},I_{s})\right)ds-\sum_{k=1}^{\infty}g_{\kappa_{k-1}\kappa_{k}}{\bf 1}_{\{\tau_{k}\leq t\}},\quad t\in[0,T].$ (5.3)
The next lemma gives the martingale characterization that lays the foundation for the loss function and the policy evaluation RL algorithm.
Lemma 5.1.
Let ${\bm{\pi}}(t,x)=(\pi_{ij}(t,x))_{i,j\in\mathbb{I}_{m}}$ be a feedback strategy and $v(t,x,i)$ be the corresponding value function given by (5.2). Then the process $M=(M_{t})_{t\in[0,T]}$ given by (5.3) is a square-integrable martingale.
Proof.
Using Itô’s rule to $v(s,X_{s},I_{s})$ from $t^{\prime}$ to $t$, we obtain
$v(t,X_{t},I_{t})$ $=v(t^{\prime},X_{t^{\prime}},I_{t^{\prime}})+\int_{t^{\prime}}^{t}(D_{x}v(s,X_{s},I_{s}))^{\top}\sigma(s,X_{s},I_{s})dW_{s}$
$\qquad+\int_{t^{\prime}}^{t}\left(\frac{\partial v(s,X_{s},I_{s})}{\partial t}+\mathcal{L}^{I_{s}}_{x}v(s,X_{s},I_{s})+\sum_{j\neq I_{s}}\left(\pi_{s}^{I_{s}j}(v(s,X_{s},j)-v(s,X_{s},I_{s}))\right)\right)ds.$ (5.4)
It follows from (5.1), (5) and (5.3) that
$\mathbb{E}[M_{t}|\mathcal{F}_{t^{\prime}}]$
$=\mathbb{E}\left[v(t,X_{t},I_{t})+\int_{0}^{t}\left(f(s,X_{s},I_{s})+\lambda R({\bm{\pi}}_{s},I_{s})\right)ds-\sum_{k=1}^{\infty}g_{\kappa_{k-1}\kappa_{k}}{\bf 1}_{\{\tau_{k}\leq t\}}\Big|\mathcal{F}_{t^{\prime}}\right]$
$=M_{t^{\prime}}+\mathbb{E}\left[v(t,X_{t},I_{t})-v(t^{\prime},X_{t^{\prime}},I_{t^{\prime}})+\int_{t^{\prime}}^{t}\left(f(s,X_{s},I_{s})+\lambda R({\bm{\pi}}_{s},I_{s})\right)ds-\int_{t^{\prime}}^{t}\sum_{j\neq I_{s}}g_{I_{s}j}\pi_{s}^{I_{s}j}ds\Big|\mathcal{F}_{t^{\prime}}\right]$
$=M_{t^{\prime}}+\mathbb{E}\left[\int_{t^{\prime}}^{t}(D_{x}v(s,X_{s},I_{s}))^{\top}\sigma(s,X_{s},I_{s})dW_{s}\Big|\mathcal{F}_{t^{\prime}}\right]=M_{t^{\prime}}.$ (5.5)
Thus, we get the desired result. ∎
Let us introduce the notation $L^{2}([0,T])$ as the space of all processes $K=(K_{t})_{t\in[0,T]}$ that $K$ is $\mathbb{F}$-progressively measurable and satisfies $\mathbb{E}[\int_{0}^{T}|K_{t}|^{2}dt]<\infty$. For any semimartingale $N=(N_{s})_{s\in[0,T]}$, we denote $L^{2}([0,T];N)$ the space of all processes $K=(K_{t})_{t\in[0,T]}$ that $K$ is $\mathbb{F}$-progressively measurable and satisfies
$\mathbb{E}\left[\int_{0}^{T}|K_{t}|^{2}d\left<N\right>_{t}\right]<\infty,$
where $\left<N\right>_{t}$ is the quadratic variation process of $N$. It follows from the martingale orthogonality condition that, for any test process $\varsigma=(\varsigma_{t})_{t\in[0,T]}\in L^{2}([0,T];M)$,
$\mathbb{E}\left[\int_{0}^{\infty}\varsigma_{t}dM_{t}\right]=0.$ (5.6)
In fact, the following result shows that this is a necessary and sufficient condition for martingale.
Proposition 5.2 (Proposition 4 in jia2022policy).
A diffusion process $N$ is a martingale if and only if
$\mathbb{E}\left[\int_{0}^{\infty}\varsigma_{t}dN_{t}\right]=0$ (5.7)
for any $\varsigma\in L^{2}([0,T];N)$.
Given a feedback strategy ${\bm{\pi}}(t,x)=(\pi_{ij}(t,x))_{i,j\in\mathbb{I}_{m}}$, we parameterize the value function using a family of functions $v^{\xi}(t,x,i)$ satisfying $v^{\xi}(T,x,i)=h(x)$, where $\xi\in\Theta\subset\mathbb{R}^{L_{\xi}}$ and $L_{\xi}$ is the dimension of the parameter vector. Let $M^{\xi}=(M_{t}^{\xi})_{t\in[0,T]}$ be the parameterized version of the martingale process $M$. Proposition 5.2 establishes that finding the optimal parameters $\xi$ reduces to solving the martingale orthogonality equation (5.7). This can be implemented through stochastic approximation with the parameter update:
$\xi\leftarrow\xi+\alpha_{\xi}\int_{0}^{T}\varsigma_{s}dM_{s}^{\xi},$ (5.8)
where $\alpha_{\xi}>0$ is the learning rate.
However, the update rule (5.8) involves a continuous-time integral that cannot be directly implemented computationally. To address this, we develop a discrete-time approximation of the martingale orthogonality condition. Let $K\in\mathbb{N}$ be the number of time intervals and $\Delta t=T/K$ be the step size. Consider the discrete partition $0=t_{0}<t_{1}<t_{2}<\cdots<t_{K}=T$ with $t_{k}-t_{k-1}=\Delta t$ for $k=1,\dots,K$. Motivated by the continuous-time update (5.8), we choose the test process $\varsigma_{t}=\frac{\partial v^{\xi}}{\partial\xi}(t,X_{t},I_{t})$ and propose the following discretized update rule to update parameters after a whole episode (offline):
$\xi\leftarrow\xi+\alpha_{\xi}\sum_{k=0}^{K-1}\frac{\partial v^{\xi}}{\partial\xi}(t_{k},X_{t_{k}},I_{t_{k}})\Delta\xi_{k}$ (5.9)
or to update parameters at every time step (online):
$\xi\leftarrow\xi+\alpha_{\xi}\frac{\partial v^{\xi}}{\partial\xi}(t_{k},X_{t_{k}},I_{t_{k}})\Delta\xi_{k}.$ (5.10)
Here $\Delta\xi_{k}$ for k=0,1,,..,K1k=0,1,,..,K-1 is given by
$\Delta\xi_{k}=v^{\xi}(t_{k+1},X_{t_{k+1}},I_{t_{k+1}})-v^{\xi}(t_{k},X_{t_{k}},I_{t_{k}})+\left(f(t_{k},X_{t_{k}},I_{t_{k}})+\lambda R({\bm{\pi}}^{\xi}_{t_{k}},I_{t_{k}})\right)\Delta t-g_{I_{t_{k}}I_{t_{k+1}}},$ (5.11)
where the parameterized strategy ${\bm{\pi}}^{\xi}(t,x)=(\pi^{\xi}_{ij}(t,x))_{i,j\in\mathbb{I}_{m}}$ is given by
$\pi_{ij}^{\xi}(t,x)=\exp\left(\frac{v^{\xi}(t,x,j)-g_{ij}-v^{\xi}(t,x,i)}{\lambda}\right),\penalty 10000\ j\neq i,$ (5.12)
and $\pi_{ii}^{\xi}(t,x)=-\sum_{j\neq i}\pi_{ij}^{\xi}(t,x)$.
Based on the above updating rules, we can present the pseudo-code of the offline PE algorithm in Algorithm 1. The online PE algorithm can be devised in a similar fashion and is omitted.
Algorithm 1 Policy Evaluation Algorithm (Offline)
Input: Initial state $(x_{0},i_{0})$, horizon $T$, number of regimes $m$, time step $\Delta t$, number of episodes $N$, number of mesh grids $K$, initial learning rates $\alpha_{\xi}(\cdot)$ (a function of the number of episodes), functional forms of parameterized value function $v^{\xi}(\cdot)$, policy $\bm{\pi}^{\xi}(\cdot)$, regime switching costs $(g_{ij})_{i,j\in\mathbb{I}_{m}}$ and temperature parameter $\lambda$.
[…]
Learning Procedure:
1:Initialize $\xi$, and $\ell=1$.
2:while $\ell<N$ do
3:  Initialize $k=0$. Observe initial state $x_{0},i_{0}$ and store $(x_{t_{0}},i_{t_{0}})\leftarrow(x_{0},i_{0})$.
4:  while $k<K$ do
5:    Generate action $j_{t_{k}}$ by $\bm{\pi}^{\xi}\left(t_{k},x_{t_{k}}\right)$.
[…]
7:    Observe new state $x$ and $i$ as output. Store $x_{t_{k+1}}\leftarrow x$, $i_{t_{k+1}}\leftarrow i$ and $f_{t_{k}}\leftarrow f$.
8:    Update $k\leftarrow k+1$.
9:  end while
10:  For every k=0,1,,K1k=0,1,...,K-1, compute
$\Delta\xi_{k}$ $=v^{\xi}(t_{k+1},x_{t_{k+1}},i_{t_{k+1}})-v^{\xi}(t_{k},x_{t_{k}},i_{t_{k}})+\left(f_{t_{k}}+\lambda R({\bm{\pi}}^{\xi}(t_{k},x_{t_{k}}),i_{t_{k}})\right)\Delta t-g_{i_{t_{k}}i_{t_{k+1}}}.$
11:  Update $\xi$ by
$\xi$ $\leftarrow\xi+\alpha_{\xi}(\ell)\sum_{k=0}^{K-1}\frac{\partial v^{\xi}}{\partial\xi}\left(t_{k},x_{t_{k}},i_{t_{k}}\right)\Delta\xi_{k},$
12:  Update $\ell\leftarrow\ell+1$.
13:end while
Proposition 4.1 and Theorem 4.2 confirm the improvement and convergence results of the policy iteration. Meanwhile, Lemma 5.1 and Proposition 5.2 show that policy evaluation can be performed by solving the martingale orthogonality condition via stochastic approximation. A natural question arises: what can be said about the convergence of Algorithm 1? To address this, we next turn to an analysis of the error estimates for Algorithm 1.
We reformulate the update rule in equation (5.9) as follows:
$\xi_{i+1}\leftarrow\xi_{i}+\alpha_{\xi}(i)\Psi(\xi_{i};X,I,{\bm{\pi}}^{\xi_{i}}),\quad i\geq 1,$ (5.13)
where
$\Psi(\xi_{i};X,I,{\bm{\pi}}^{\xi_{i}})=\sum_{k=0}^{K-1}\frac{\partial v^{\xi}}{\partial\xi}(t_{k},X_{t_{k}},I_{t_{k}})\Delta\xi_{k},$
with $\Delta\xi_{k}$ defined in equation (5.11). For notational convenience, we introduce the shorthand $Y_{i+1}=(X,I,{\bm{\pi}}^{\xi_{i}})$ for $i\geq 1$. We further define the expected update function as $\psi(\xi):=\mathbb{E}[\Psi(\xi;Y)]$. To establish convergence guarantees, we make the following technical assumptions.
Assumption 5.3.
  • (i)
    The ordinary differential equation $\xi^{\prime}(t)=\psi(\xi(t))$ has a unique stable equilibrium point $\xi^{*}$.
  • (ii)
    There exists a constant $C>0$ such that $\mathbb{E}[|\Psi(\xi_{i};Y_{i+1})|^{2}|\xi_{i}]\leq C(1+|\xi_{i}|^{2})$ for all iterations.
  • (iii)
    There exists $\kappa>0$ such that $(\xi-\xi^{*})\cdot\psi(\xi)\leq-\kappa|\xi-\xi^{*}|^{2}$ for all $\xi\in\mathbb{R}^{L_{\xi}}$.
  • (iv)
    There exist constants $\rho,C>0$ such that $\sup_{j\in\mathbb{I}_{m}}|v^{\xi}(\cdot,j)-v^{\xi^{*}}(\cdot,j)|_{C^{0}(\overline{{\cal D}})}\leq C|\xi-\xi^{*}|^{\rho}$ for all $\xi\in\mathbb{R}^{L_{\xi}}$.
Under these conditions, we now present the main convergence result, which provides the explicit error bound for Algorithm 1.
Theorem 5.4.
Let Assumption 5.3 hold. Set $\alpha_{\xi}(i)=\frac{A}{i^{\nu}+B}$ for some $\nu\leq 1$, $A>\frac{\nu}{2\kappa}$ and $B>0$, and let $\epsilon>0$. Then there exists $C>0$ (independent of $n,\epsilon$) such that with probability of at least $1-\epsilon$,
$\sup_{j\in\mathbb{I}_{m}}|v^{\xi_{i}}(\cdot,j)-v(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}\leq\sup_{j\in\mathbb{I}_{m}}|v(\cdot,j)-v^{\xi^{*}}(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}+\frac{C}{\epsilon^{\rho/2}}i^{-\frac{\nu\rho}{2}}.$ (5.14)
Proof.
Under Assumptions 5.3 (i)–(iii) and the step-size condition on $\alpha_{\xi}(i)$, an application of Theorem 22 in benveniste2012adaptive yields
$\mathbb{E}[|\xi_{i}-\xi^{*}|^{2}]\leq Ci^{-\nu}$
where $C>0$ is a constant independent of $n$. This bound in turn implies that
$|\xi_{i}-\xi^{*}|\leq C\epsilon^{-\frac{1}{2}}i^{-\frac{\nu}{2}}$
with probability at least $1-\epsilon$. Then, invoking Assumption 5.3 (iv), we deduce that with probability at least $1-\epsilon$,
$\sup_{j\in\mathbb{I}_{m}}|v^{\xi_{i}}(\cdot,j)-v(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}$
$\leq\sup_{j\in\mathbb{I}_{m}}|v(\cdot,j)-v^{\xi^{*}}(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}+\sup_{j\in\mathbb{I}_{m}}|v^{\xi^{*}}(\cdot,j)-v^{\xi_{i}}(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}$
$\leq\sup_{j\in\mathbb{I}_{m}}|v^{\xi^{*}}(\cdot,)-v(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}+\frac{C}{\epsilon^{\rho/2}}i^{-\frac{\nu\rho}{2}}.$
This completes the proof of the theorem. ∎
Theorem 5.4 establishes a comprehensive error analysis for Algorithm 1, providing both theoretical guarantees and practical insights into its convergence behavior. The result demonstrates that the policy evaluation error can be systematically decomposed into two distinct components: the approximation error of the parametric function class and the algorithmic error arising from the stochastic approximation procedure. The first term, $\sup_{j\in\mathbb{I}_{m}}|v(\cdot,j)-v^{\xi^{*}}(\cdot,j)|_{C^{0}({\overline{{\cal D}}})}$, represents the inherent approximation capability of our chosen parametric family. This bias term is independent of the learning algorithm and reflects how well the optimal parameter $\xi^{*}$ can approximate the true value function within the selected function class. The second term, $Ci^{-\frac{\nu\rho}{2}}/\epsilon^{\rho/2}$, exhibits a polynomial decay with respect to the iteration number $i$ and vanishes asymptotically as the number of iterations increases, demonstrating the algorithm’s convergence to the optimal parameter configuration within the chosen function class.
6 Numerical Examples
This section presents some numerical experiments to demonstrate the practical efficacy of the proposed RL algorithm. We first examine a bounded regulator problem to analyze the algorithm’s convergence property and policy behavior. Subsequently, we apply the algorithm to a put option selection problem involving the optimal switching between risky assets, showcasing its effectiveness in a more complex, multi-regime setting with some financial interpretations.
6.1 Bounded Regulator Problem
To establish a performance benchmark for our algorithm, we consider a finite-horizon optimal switching problem with two regimes, conceptualized as a bounded regulator. This classic problem provides a tractable yet non-trivial testbed where the optimal policy has an intuitive structure, allowing for clear interpretation of the algorithm’s learned strategy.
The system state $X=(X_{t})_{t\in[0,T]}$ evolves according to regime-specific stochastic dynamics:
$dX_{t}=\mu_{i}dt+\sigma dW_{t},\quad i\in\{0,1\},\penalty 10000\ t\in[0,T],$ (6.1)
with initial condition $X_{0}=x\in\mathbb{R}$. Here, $W=(W_{t})_{t\in[0,T]}$ is a standard Brownian motion. The parameters are chosen with symmetry: the drift coefficients are $\mu_{0}=-2$ and $\mu_{1}=2$, and the volatility is $\sigma=0.5$. This symmetric setup induces a natural switching logic to correct the state’s deviation.
The controller’s objective is to maximize the expected total reward over the horizon $[0,T]$, which comprises a running reward and a terminal reward:
$f(x)=2e^{-2x^{2}}-0.1,\quad h(x)=2e^{-2x^{2}},\quad x\in\mathds{R}.$
The Gaussian bump shape of the functions $f$ and $h$ creates a strong incentive to maintain the state $X_{t}$ near zero, as the reward attains its maximum value at $x=0$. Each switch between regimes incurs a cost, specified as $g_{01}=g_{10}=0.5$. This cost penalizes excessive control actions, forcing the optimal policy to strategically balance the benefit of corrective switching against the incurred cost.
We use a discrete version of (6.1) for $t=0,\Delta t,...,K\Delta t$ with $T=1$, $K=100$ and $\Delta t=T/K$. The value function and policy are approximated by a neural network in the PyTorch framework with the architecture and parameters summarized in Table 1.
In order to approximate the expectation operator in the theoretical formulation via Monte Carlo methods—that is, to estimate the gradient direction using the average of finite samples so that the update of parameters $\xi$ is not based on a single specific scenario—we employ batch learning to train the neural network parameters $\xi$. This approach not only enhances training stability but also ensures that the learned policy is optimal in an expected sense.
Table 1: Neural Network Architecture and Training Parameters for the Regulator Problem
Component Specification
Network Architecture 2 hidden layers
Activation Functions ReLU (Layer 1), Tanh (Layer 2)
Hidden Dimension 128
Batch Size 64
Optimizer Adam
Learning Rate $1\times 10^{-3}$
Training Episodes 1000
The training progression under the temperature parameter $\lambda=0.2$ is shown in Figure 1-(a). The loss function decreases efficiently and stabilizes after approximately 400 episodes, indicating the robust convergence of the policy iteration in the RL algorithm. Figure 1-(b) depicts the learned value functions and the corresponding switching probabilities at $t=0.5$. The near symmetry between the value functions for regime 0 (blue line) and regime 1 (orange line) is a direct consequence of the symmetric problem parameters. The switching probabilities—from regime 0 to 1 (green line) and from regime 1 to 0 (yellow line)—are calculated from the optimal intensity $\pi$.
Figure 1: (a): Convergence of the training loss for the bounded regulator problem with $\lambda=0.2$. (b): Learned value functions and switching probabilities at $t=0.5$ and $I_{0.5}=0$ for $\lambda=0.2$, i.e, $P(I_{t+\Delta t}=1|I_{t}=0,X_{t}=x,t=0.5)$.
A central theoretical result is the convergence of the exploratory solution to the classical optimal switching policy as the temperature parameter $\lambda$ tends to zero. We validate this numerically. Figure 2 shows that the training loss decreases for different values of $\lambda$, with convergence achieved in all cases. More importantly, Figure 3 illustrates the fundamental transformation of the optimal policy. For a larger $\lambda$ (e.g., 0.2), the switching probability is a smooth function of the state, reflecting exploratory randomization. As $\lambda$ decreases to 0.01, the probability curve becomes sharp and nearly binary, approaching a deterministic threshold-based policy. This visual evidence strongly supports the theoretical finding that the solutions of the exploratory HJB equations converge to the solution of the classical variational inequalities as $\lambda\to 0$.
Figure 2: Training convergence for different temperature parameters $\lambda$.
Figure 3: Evolution of the switching probability from regime 0 to 1 as $\lambda$ decreases.
6.2 Put Option Selection Problem
To demonstrate the algorithm’s applicability in finance, we model an investor who aims to optimally switch an investment decision between three regimes: two European put options on different assets and a risk-free savings account. The investor’s wealth can be allocated to one of three regimes during the finite horizon $[0,T]$:
  • regime 0: a put option on Stock $A$.
  • regime 1: a put option on Stock $B$.
  • regime 2: the risk-free savings account.
The underlying stock prices follow the geometric Brownian motion:
$dS^{A}_{t}=\mu^{A}S^{A}_{t}dt+\sigma^{A}S^{A}_{t}dW_{t},\quad dS^{B}_{t}=\mu^{B}S^{B}_{t}dt+\sigma^{B}S^{B}_{t}dW_{t},\quad t\in(0,T],$
with $S^{A}_{0}=s^{A}\in[0,\infty),S^{B}_{0}=s^{B}\in[0,\infty)$. Here the parameters are set by $(\mu^{A},\sigma^{A})=(0.1,0.2)$ and $(\mu^{B},\sigma^{B})=(0.05,0.1)$, and $W=(W_{t})_{t\in[0,T]}$ is a standard Brownian motion. The risk free rate is $r=0.05$. For any time $t\in[0,T]$, the investor decides an action $I_{t}\in\{0,1,2\}$, which determines the regime in which the investor’s wealth is allocated. Switching between regimes incurs transaction costs given by the matrix:
$G=(g_{ij})_{0\leq i,j\leq 2}=\begin{bmatrix}0&0.02&0.01\\ 0.02&0&0.01\\ 0.02&0.02&0\end{bmatrix}$.
The investor’s objective is to maximize the expected total reward over the horizon $[0,T]$, where the running reward function is given by
$f(s^{A},s^{B},i)=\begin{cases}(S_{K}-s^{A})^{+},&i=0,\\ (S_{K}-s^{B})^{+},&i=1,\\ rS_{K},&i=2,\end{cases}$
with the strike price $S_{K}=1$ and $(x)^{+}:=\max\{x,0\}$ for $x\in\mathds{R}$. The terminal reward function is assumed to be 0.
We set the time horizon $T=1$, the number of time intervals $K=50$, the step size $\Delta t=T/K=0.02$, and the temperature parameter $\lambda=0.1$. The value function and policy are approximated by a neural network with the architecture and parameters summarized in Table 2. The model was implemented within the PyTorch framework.
Table 2: Neural Network Architecture and Training Parameters for the Regulator Problem
Component Specification
Network Architecture 2 hidden layers
Activation Functions Tanh (Layer 1), Tanh (Layer 2)
Hidden Dimension 128
Batch Size 512
Optimizer Adam
Learning Rate $1\times 10^{-4}$
Training Episodes 1000
According to Figure 4, at the beginning of training, the loss exhibits a oscillation, and the convergence is very pronounced. It becomes stable when the number of episodes exceeds 800. We then plot in Figure 5 the learnt value functions $(V^{1},V^{2},V^{3})$. Conforming to financial intuition, each value function $V^{i}$ (for $i=1,2,3$) is a decreasing function of both underlying stock prices $s^{A}$ and $s^{B}$.
Figure 4: The training loss for the put option selection problem.
Figure 5: Left panel: The learnt value functions $s^{A}\to(V^{1},V^{2},V^{3})$ with $(s^{B},t)=(1.0,0.5)$. Right panel: The learnt value functions $s^{B}\to(V^{1},V^{2},V^{3})$ with $(s^{A},t)=(1.0,0.5)$.
Acknowledgements: Yijie Huang, Mengge Li and Xiang Yu are supported by the Hong Kong RGC General Research Fund (GRF) under grant no. 15211524 and the Hong Kong Polytechnic University research grant under no. P0045654.